Hagy Corporation has an activity-based costing system with three activity cost pools--Processing, Setting Up, and Other. The company's overhead costs, which consist of factory utilities and indirect labor, are allocated to the cost pools in proportion to the activity cost pools' consumption of resources. Data concerning the company's costs and activity-based costing system appear below:

Factory utilities (total) $ 107,000
Indirect labor (total) $ 9,500

Distribution of Resource Consumption Across Activity Cost Pools

Processing Setting Up Other
Factory utilities 0.10 0.50 0.40
Indirect labor 0.10 0.20 0.70
Activity Cost Pools
Processing Setting Up Other Total
Factory utilities $0
Indirect labor $0
Total $0 $0 $0 $0

Required:

Assign overhead costs to activity cost pools using activity-based costing.

Activity Cost Pools
Processing Setting Up Other Total
Factory utilities
Indirect labor
Total
Answer #1
Activity Cost Pools
Processing Setting Up Other Total
Factory utilities 10700 53500 42800 107000
Indirect labor 950 1900 6650 9500
Total 11650 55400 49450 116500
Workings:
Processing Setting Up Other
Factory utilities =107000*0.10 =107000*0.50 =107000*0.40
Indirect labor =9500*0.10 =9500*0.20 =9500*0.70